OpenStudy (anonymous):

The Diaz family is driving at a constant speed on the highway so their distance varies directly with their speed. They traveled 17.5 miles in 15 minutes. How far did they travel in 2 hours?

OpenStudy (bahrom7893):

17.5 - 15 x - 120 x = 120*17.5/15 = 140

OpenStudy (bahrom7893):

there are 120 mins in one hour.

OpenStudy (bahrom7893):

i meant in 2 hrs

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Do you have any question?

OpenStudy (anonymous):

Nope :3